# What Comes After Harm? Mapping Reparative Actions in AI through Justice Frameworks

### v1

@@ Introduction
When New York City introduced a law requiring employers to disclose their use of AI systems in hiring and to publish audit results (New York City Department of Consumer and Worker Protection 2021), it was widely celebrated as a breakthrough in algorithmic accountability. Policymakers and researchers saw it as a promising intervention to address bias and discrimination and to drive organizational change (Johnson 2021). Yet, the law stops short of requiring companies to act on audit findings, such as correcting disparate impact. Critics have pointed out that organizations can comply in form but not in substance by disclosing superficial results or by inserting a human reviewer— workarounds that create the appearance of accountability without delivering meaningful reform (Wright et al. 2024).
This example illustrates a gap between recognizing harm and taking responsibility for addressing it. While recent research in Responsible AI (RAI) has made substantial progress in identifying and characterizing harms (Schelble et al. 2024; Zhang et al. 2024; Techtonic Justice 2023) and in enhancing the auditing process (Mökander 2023; Metaxa et al. 2021), we still know relatively little about what actions are taken after harm is acknowledged. In particular, it remains unclear how often those actions contribute to meaningful redress for those impacted, which is critical for moving beyond the appearance of oversight to meaningful accountability.
In this paper, we focus on the notion of reparative action. Grounded in theories of justice, we view reparation not merely as a pragmatic fix, but as a normative response to wrongdoing (United Nations General Assembly 2005). We draw from three justice frameworks widely applied in the context of social and technological harm: punitive, restorative, and transformative justice (Feinberg 2019; Zehr 2015; Nocella and Anthony 2011). Using these frameworks, we define “AI harm reparative action” as the process of remedying the harm caused by AI systems and restoring justice for those affected.
@@
We draw on theories of justice to interpret what meaningful reparation could look like. Punitive justice emphasizes holding perpetrators accountable through consequences; restorative justice focuses on the needs and voices of those harmed; and transformative justice seeks to address the structural conditions that enabled harm. Yet current practices often fall short of these ideals, lacking clear accountability, offering limited support to affected communities, and rarely confronting systemic causes. We also highlight the vital but often overlooked role of affected communities and civil society in initiating responses, and suggest how future reparation practices might engage them more fully.
@@
We follow a normative definition of AI harm as a setback to an individual or group’s interests caused by the design, deployment, or use of an AI system (Diberardino, Baleshta, and Stark 2024). Existing research largely emphasizes harm detection at the design and development stages and evaluation after AI is created  (Raji et al. 2020). Researchers and practitioners increasingly turn to AI auditing, which involves systematic evaluations of AI systems to assess compliance with ethical, legal, or organizational standards (Mökander 2023; Metaxa et al. 2021; Birhane et al. 2024). A related technique is AI red-teaming which adopts a more adversarial perspective on stress-testing AI systems and simulation of bad actors or harmful use cases to reveal emergent or latent risks (Ganguli et al. 2022). On the other hand, benchmarking tools such as UnsafeBench (Qu et al. 2024) and AgentHarm (Andriushchenko et al. 2024) enable quantification of potential risks.
While these approaches are increasingly applied to post-deployment systems (Casper et al. 2024; Singh et al. 2025), the goal remains oriented toward harm prevention and risk detection. Few provide mechanisms to address harm once it has occurred or to involve affected parties in shaping the response. To address this gap, emerging research has proposed participatory methods that incorporate users into the auditing and red-teaming process (Lam et al. 2022; Cabrera et al. 2021; Shen et al. 2021; Deng et al. 2023, 2025). These approaches draw on users’ situated knowledge to uncover harms that technical evaluations may overlook. However, scholars have questioned whether these methods truly foster accountability or merely serve as corporate risk management strategies (Gillespie et al. 2024; Feffer et al. 2024). Our work complements this line of work on detection and evaluation by shifting focus to the aftermath of harm. We examine how AI harm is acknowledged and repaired: who takes action, in what form, and whether these responses align with the needs of those affected.
@@
While much attention has focused on preventing AI harm, a growing body of research examines how individuals and communities respond after harm has occurred. Studies have documented acts of resistance and refusal by affected users and workers (Velkova and Kaun 2021; Ganesh and Moss 2022), as well as conditions that prompt these responses. For instance, Johnson et al. analyzed the abandonment of algorithms following public backlash (Johnson et al. 2024), while Ehsan et al. show how withdrawn systems can still erode institutional trust (Ehsan et al. 2022). Others have explored how people believe AI harms should be addressed, such as public preferences for punishment or accountability (Lima, Grgić-Hlača, and Cha 2023).
Together, these studies shed light on diverse post-harm responses, yet most focus on isolated cases, specific stakeholder groups, or relatively narrow accountability mechanisms. As a result, we lack a system-level understanding of how different actors respond after harm is surfaced and whether these responses reflect principles of accountability or reparation (Bogiatzis-Gibbons 2024). DeVrio et al.’s taxonomy offers a foundational framework by categorizing grassroots responses to algorithmic harm, emphasizing the roles of power dynamics in these interactions (DeVrio, Eslami, and Holstein 2024). Our study builds on this work by analyzing over 1,000 AI harm incidents to map stakeholder actions at scale and identify where reparative efforts occur or are absent. By focusing on reparation, we move beyond harm detection to examine the obligations that arise once harm becomes visible.
@@
This paper draws on three major justice traditions—punitive, restorative, and transformative justice—to contextualize different approaches to reparation. Punitive justice emphasizes holding perpetrators accountable through proportional punishment, such as incarceration, fines, or legal sanctions (Feinberg 2019; Foucault 2023; Ashworth and Kelly 2021). It is the dominant model in most legal systems and widely used across online and offline governance (Gillespie 2018). In contrast, restorative justice centers the needs of those harmed, aiming to repair relationships and rebuild trust through practices such as public acknowledgment, apology, dialogue, and community-based support (Zehr 2015; Pranis 2015). It has been widely applied in criminal justice and increasingly studied in digital governance and online harm contexts (Xiao, Jhaver, and Salehi 2023; Xiao et al. 2020). Transformative justice builds on these traditions by addressing the structural conditions that enable harm, such as racism, ableism, and economic injustice (Morris 2000; Kaba 2021). It expands the idea of reparation beyond individual or relational repair, advocating for systemic change through institutional reform, power redistribution, and community-led design (Kaba and Hassan 2019).
These frameworks offer complementary perspectives in how they define harm, assign responsibility, and envision repair. Punitive and restorative justice both focus on interpersonal accountability, but take different approaches: punitive justice emphasizes external consequences such as punishment and deterrence, while restorative justice emphasizes mutual recognition, healing, and support for those affected (Wenzel et al. 2008). In addition, while restorative and transformative justice both center the harmed party, they differ in whether the aim is to restore relationships or to transform institutions (Kim 2021).
These theories of justice have been widely applied to discussions of harm reparation in offline contexts and the governance of online communities (Nocella and Anthony 2011; Chordia et al. 2024). In the context of harm caused by AI systems, however, the process of reparation remains underexplored, and little attention has been given to the normative frameworks needed to assess the adequacy of such responses. In this study, we use these justice frameworks not only to interpret reparative actions, but also to evaluate how these actions contribute to meaningful repair.
@@
We based our analysis on the AIAAIC repository of AI incidents, an independent, public-interest resource that documents events and controversies involving AI, algorithms, and automation [^2] https://www.aiaaic.org/. Compared to similar resources (e.g., the AI Incident Database), AIAAIC offers greater breadth and detail, with over 1000 incidents that include summaries, analyses of causes and implications, and links to related news coverage. It has been maintained by global wide volunteers and has been widely used in AI harm research to examine the consequences of AI failures and to develop harm taxonomies (Lee et al. 2024a; Johnson et al. 2024).
@@
Using this criterion, we applied purposive sampling (Campbell et al. 2020) to identify 671 relevant incidents. From these, we randomly selected 200 to form the taxonomy development dataset. To ensure analytical saturation, we spot-coded 20 additional incidents (5% of the remainder) (Lee et al. 2024b). No new action types emerged, indicating that the sample sufficiently captured the range of reparative responses.
@@
To construct our taxonomy of AI harm reparation, we conducted a thematic analysis (Creswell and Creswell 2017) of the 200-incident taxonomy development dataset. While our analysis was primarily inductive, it was informed by concepts and language drawn from restorative, punitive, and transformative justice, which offered conceptual grounding for distinguishing types of reparation.
@@
We next examine how these reparative actions are distributed in the AIAAIC incident database. To address this, we applied the taxonomy developed in the previous stage to all 1060 incidents. Given the scale of the dataset, manual coding was impractical. Recent research has demonstrated the potential of large language models (LLMs) to support deductive qualitative coding at scale (Xiao et al. 2023; Mun et al. 2024). We used GPT-4 Turbo, accessed via the OpenAI API [^3] https://platform.openai.com, to assist with this multi-label classification task, in which the model identified the presence or absence of each predefined reparative action and stakeholder category within incident descriptions.
We scraped each incident’s summary and linked articles to form a text corpus, then applied a two-step prompting pipeline. First, the model extracted stakeholder-action pairs, mirroring the structure of our manual coding. Second, it assessed whether each action category was present, classified responsible stakeholders, and provided supporting evidence. An example prompt is included in Appendix C. To validate this approach, we compared model outputs to human annotations for a random sample of 200 incidents (20 per action category). Three coders labeled action and stakeholder categories, achieving 93% inter-coder reliability for actions and 90% for stakeholders. Using these annotations as ground truth, the LLM achieved 87% accuracy on actions and 79% on stakeholders, indicating substantial agreement. Based on this validation, we applied the LLM to code the remaining incidents. Summary statistics and prompt details are provided in Appendix D.
@@
Our analysis draws on the AIAAIC repository, which aggregates incidents primarily from public sources such as news articles. As a result, it may overrepresent high-profile cases while overlooking less visible harms. Some stakeholder actions—such as private responses by affected individuals (DeVrio, Eslami, and Holstein 2024) or undisclosed internal changes by organizations (Lukpat )—may be difficult to detect, potentially leading to underestimation of involvement and remedy.
To scale analysis across the dataset, we used large language models. While we validated model outputs against human-coded samples and carefully designed prompts, LLM performance remains shaped by training data and prompt construction (Guo et al. 2024). We treat LLM-assisted coding as a tool for identifying large-scale patterns, not as a replacement for human interpretation.
@@ A Taxonomy of AI Harm Reparative Actions
We present a taxonomy of AI harm reparative actions (Table 1), organized according to four overarching goals that reflect key orientations in justice theory: Acknowledgment, Attribution, Remedy, and Reform. Rather than assigning each action to a specific justice framework, we used concepts from punitive, restorative, and transformative justice to group actions based on their underlying reparative aims. Acknowledgment and Attribution are foundational steps shared across justice traditions: addressing harm begins with recognizing its occurrence and establishing responsibility (United Nations General Assembly 2005). Remedy aligns with both punitive and restorative justice, though its emphasis is on redress and interpersonal accountability (Zehr 2015; Duff 2005). Reform, on the other hand, reflects the commitments of transformative justice, which seeks to address systemic conditions and prevent future harm (Morris 2000).
@@
To contextualize the taxonomy and illustrate broader patterns, we present a high-level summary of how frequently each reparative goal appears across the dataset. While the taxonomy is qualitatively developed to reflect justice-oriented goals and actions, this accompanying quantitative overview provides an empirical lens for understanding the prevalence and distribution of reparative responses in practice.
In our analysis of 1,060 AI incidents, 54% involved actions related to Acknowledgment and 47% involved Attribution. By contrast, only 15% of cases included any form of Remedy, and 28% showed evidence of Reform. Figure 1 presents the distribution of specific actions across these goals. The most frequent action was communication from perpetrators, appearing in 541 cases (51%). In comparison, more impactful forms of follow-through were rare: only 7% of cases involved legal or policy reform, 6% mentioned compensation, and just 4% resulted in the discontinuation of a product or feature.
@@
Attribution Investigation of Harm Regulators, Media, AI Corporations Internal corporate audits; Government audits; Investigative journalism
Initiation of Lawsuits Affected Users, Advocacy Groups, Regulators Civil or regulatory lawsuits
Remedy Repercussions Regulators, AI Corporations Fines; Criminal charges; Dismissals
Compensation AI Corporations, End-Users Refunds; Settlements
Reform AI Design Changes AI Corporations Algorithm redesign; Adjustments to product safety protocols
Content or Product Removal AI Corporations Content takedown; Product recalls
Discontinuation AI Corporations Shutting down chatbots or features
Law or Policy Change Regulators, AI Corporations New laws; Internal policy updates
Table 1: Taxonomy of Reparative Actions. This table presents our taxonomy of reparative actions, organized by four goals: Acknowledgment, Attribution, Remedy, and Reform. Each action lists the initiating stakeholders and examples. The taxonomy highlights the range of justice orientations and multi-stakeholder participation in AI harm responses.

### v2

@@ Introduction
When New York City introduced a law requiring employers to disclose their use of AI systems in hiring and to publish audit results (nyc_aedt_2021), it was widely celebrated as a breakthrough in algorithmic accountability. Policymakers and researchers saw it as a promising intervention to address bias and discrimination and to drive organizational change (johnson_ai_accountability_2021). Yet, the law stops short of requiring companies to act on audit findings, such as correcting disparate impact. Critics have pointed out that organizations can comply in form but not in substance by disclosing superficial results or by inserting a human reviewer— workarounds that create the appearance of accountability without delivering meaningful reform (wright2024null).
This example illustrates a gap between recognizing harm and taking responsibility for addressing it. While recent research in Responsible AI (RAI) has made substantial progress in identifying and characterizing harms (schelble_towards_2024; zhang2024dark; techtonic2023inescapable) and in enhancing the auditing process (mokander2023auditing; metaxa_auditing_2021), we still know relatively little about what actions are taken after harm is acknowledged. In particular, it remains unclear how often those actions contribute to meaningful redress for those impacted, which is critical for moving beyond the appearance of oversight to meaningful accountability.
In this paper, we focus on the notion of reparative action. Grounded in theories of justice, we view reparation not merely as a pragmatic fix, but as a normative response to wrongdoing (un_reparation_2005). We draw from three justice frameworks widely applied in the context of social and technological harm: punitive, restorative, and transformative justice (feinberg2019expressive; zehr_little_2015; nocella2011overview). Using these frameworks, we define “AI harm reparative action” as the action taken in response to harm caused by AI systems that participates in the negotiation of redress and justice. In this view, reparative action is not a single event, but a range of practices that play a role in shaping how accountability and repair are understood and enacted.
@@
We draw on theories of justice to interpret what meaningful reparation could look like. Punitive justice emphasizes holding perpetrators accountable through consequences; restorative justice focuses on the needs and voices of those harmed; and transformative justice seeks to address the structural conditions that enabled harm. Yet current practices often fall short of these ideals, lacking clear accountability, offering limited support to affected communities, and rarely confronting systemic causes. We also highlight the vital but often overlooked role of affected communities and civil society in initiating responses, and suggest how future reparation practices might engage them.  [^2] A full version of this paper (with appendix) is at https://arxiv.org/abs/2506.05687.
@@
We follow a normative definition of AI harm as a setback to an individual or group’s interests caused by the design, deployment, or use of an AI system (diberardino2024algorithmic). Existing research largely emphasizes harm detection at the design and development stages and evaluation after AI is created  (raji2020closing). Researchers and practitioners increasingly turn to AI auditing, which involves systematic evaluations of AI systems to assess compliance with ethical, legal, or organizational standards (mokander2023auditing; metaxa_auditing_2021; birhane2024ai). A related technique is AI red-teaming which adopts a more adversarial perspective on stress-testing AI systems and simulation of bad actors or harmful use cases to reveal emergent or latent risks (ganguli2022red). On the other hand, benchmarking tools such as UnsafeBench (qu2024unsafebench) and AgentHarm (andriushchenko2024agentharm) enable quantification of potential risks.
While these approaches are increasingly applied to post-deployment systems (casper2024black; singh2025red), the goal remains oriented toward harm prevention and risk detection. Few provide mechanisms to address harm once it has occurred or to involve affected parties in shaping the response. To address this gap, emerging research has proposed participatory methods that incorporate users into the auditing and red-teaming process (lam_end-user_2022; cabrera2021discovering; shen_everyday_2021; deng_understanding_2023; deng2025weaudit). These approaches draw on users’ situated knowledge to uncover harms that technical evaluations may overlook. However, scholars have questioned whether these methods truly foster accountability or merely serve as corporate risk management strategies (gillespie2024ai; feffer_red-teaming_2024). Our work complements this line of work on detection and evaluation by shifting focus to the aftermath of harm. We examine how AI harm is acknowledged and repaired: who takes action, in what form, and whether these responses align with the needs of those affected.
@@
While much attention has focused on preventing AI harm, a growing body of research examines how individuals and communities respond after harm has occurred. Studies have documented acts of resistance and refusal by affected users and workers (velkova_algorithmic_2021; ganesh2022resistance), as well as conditions that prompt these responses. For instance, Johnson et al. analyzed the abandonment of algorithms following public backlash (johnson_fall_2024), while Ehsan et al. show how withdrawn systems can still erode institutional trust (ehsan_algorithmic_2022). Others have explored how people believe AI harms should be addressed, such as public preferences for punishment or accountability (lima_blaming_2023), or how grassroots actors organize for redress in the face of institutional power (devrio_building_2024).
This growing literature has shed light on varied responses to harm, but often centers on individual cases, specific actors, or limited forms of accountability (bogiatzis2024beyond). Our study contributes a systematic analysis of over 1,000 AI harm incidents. Rather than treating responses as isolated events, we conceptualize reparation as a series of actions unfolding over time and involving multiple stakeholders. In addition, we interpret these responses through the lens of reparation and justice, examining how obligations to repair harm are acknowledged, resisted, or enacted in practice.
@@
This paper draws on three major justice traditions—punitive, restorative, and transformative justice—to contextualize different approaches to reparation. Punitive justice emphasizes holding perpetrators accountable through proportional punishment, such as incarceration, fines, or legal sanctions (feinberg2019expressive; foucault2023discipline; ashworth2021sentencing). It is the dominant model in most legal systems and widely used across online and offline governance (gillespie_custodians_2018). In contrast, restorative justice centers the needs of those harmed, aiming to repair relationships and rebuild trust through practices such as public acknowledgment, apology, dialogue, and community-based support (zehr_little_2015; pranis_little_2015). It has been widely applied in criminal justice and increasingly studied in digital governance and online harm contexts (xiao_addressing_2023; xiao_random_2020). Transformative justice builds on these traditions by addressing the structural conditions that enable harm, such as racism, ableism, and economic injustice (morris2000stories; kaba_we_2021). It expands the idea of reparation beyond individual or relational repair, advocating for systemic change through institutional reform, power redistribution, and community-led design (kaba_fumbling_2019).
These frameworks offer complementary perspectives in how they define harm, assign responsibility, and envision repair. Punitive and restorative justice both focus on interpersonal accountability, but take different approaches: punitive justice emphasizes external consequences such as punishment and deterrence, while restorative justice emphasizes mutual recognition, healing, and support for those affected (wenzel2008retributive). In addition, while restorative and transformative justice both center the harmed party, they differ in whether the aim is to restore relationships or to transform institutions (kim2021transformative).
These theories of justice have been widely applied to discussions of harm reparation in offline contexts and the governance of online communities (nocella2011overview; 10.1145/3613904.3642704). In the context of harm caused by AI systems, however, the process of reparation remains underexplored, and little attention has been given to the normative frameworks needed to assess the adequacy of such responses. In this study, we use these justice frameworks not only to interpret reparative actions, but also to evaluate how these actions contribute to meaningful repair.
@@
We based our analysis on the AIAAIC repository of AI incidents, an independent, public-interest resource that documents events and controversies involving AI, algorithms, and automation [^3] https://www.aiaaic.org/. Compared to similar resources (e.g., the AI Incident Database), AIAAIC offers greater breadth and detail, with over 1000 incidents that include summaries, analyses of causes and implications, and links to related news coverage. It has been maintained by global wide volunteers and has been widely used in AI harm research to examine the consequences of AI failures and to develop harm taxonomies (lee2024deepfakes; johnson_fall_2024).
@@
Using this criterion, we applied purposive sampling (campbell2020purposive) to identify 671 relevant incidents. From these, we randomly selected 200 to form the taxonomy development dataset. To ensure analytical saturation, we spot-coded 20 additional incidents (5% of the remainder) (lee_deepfakes_2024). No new action types emerged, indicating that the sample sufficiently captured the range of reparative responses.
@@
To construct our taxonomy of AI harm reparation, we conducted a thematic analysis (creswell_research_2017) of the 200-incident taxonomy development dataset. While our analysis was primarily inductive, it was informed by concepts and language drawn from restorative, punitive, and transformative justice, which offered conceptual grounding for distinguishing types of reparation.
@@
We next examine how these reparative actions are distributed in the AIAAIC incident database. To address this, we applied the taxonomy developed in the previous stage to all 1060 incidents. Given the scale of the dataset, manual coding was impractical. Recent research has demonstrated the potential of large language models (LLMs) to support deductive qualitative coding at scale (xiao2023supporting; mun2024particip). We used GPT-4 Turbo, accessed via the OpenAI API [^4] https://platform.openai.com, to assist with this multi-label classification task, in which the model identified the presence or absence of each predefined reparative action and stakeholder category within incident descriptions.
We scraped each incident’s summary and linked articles to form a text corpus, then applied a two-step prompting pipeline. First, the model extracted stakeholder-action pairs, mirroring the structure of our manual coding. Second, it assessed whether each action category was present, classified responsible stakeholders, and provided supporting evidence. During prompt development, we conducted multiple rounds of qualitative error analysis to identify common failure patterns and refine the prompt. An example of the final prompt is included in Appendix C. To validate this approach, we compared model outputs to human annotations for a random sample of 200 incidents (20 per action category). Three coders labeled action and stakeholder categories, achieving 93% inter-coder reliability for actions and 90% for stakeholders. Using these annotations as ground truth, the LLM achieved 87% accuracy on actions and 79% on stakeholders, indicating substantial agreement. Based on this validation, we applied the LLM to code the remaining incidents. Summary statistics and prompt details are provided in Appendix D.
@@
Our analysis draws on the AIAAIC repository, which aggregates incidents primarily from public sources such as news articles. As a result, the dataset is not a comprehensive or random sample of AI harm and may overrepresent high-profile cases while overlooking less visible or underreported incidents. Certain responses—such as private actions by affected individuals (devrio_building_2024) or undisclosed internal changes by organizations (lukpat_ai_employees_2024)—are less likely to appear in public reporting, potentially leading to an underestimation of some stakeholders’ involvement.
To scale our analysis, we used large language models (LLMs). While we validated outputs against human-coded samples and carefully designed prompts, LLM performance is shaped by training data and prompt design (guo2024evaluating). We treat LLM-assisted coding as a tool to identify broad patterns, not a substitute for human interpretation.
@@ A Taxonomy of AI Harm Reparative Actions
We present a taxonomy of AI harm reparative actions (Table 1), organized according to four overarching goals that reflect key orientations in justice theory: Acknowledgment, Attribution, Remedy, and Reform. Rather than assigning each action to a specific justice framework, we used concepts from punitive, restorative, and transformative justice to group actions based on their underlying reparative aims. Acknowledgment and Attribution are foundational steps shared across justice traditions: addressing harm begins with recognizing its occurrence and establishing responsibility (un_reparation_2005). Remedy aligns with both punitive and restorative justice, though its emphasis is on redress and interpersonal accountability (zehr_little_2015; duff2005punishment). Reform, on the other hand, reflects the commitments of transformative justice, which seeks to address systemic conditions and prevent future harm (morris2000stories).
@@
To contextualize the taxonomy and illustrate broader patterns, we provide a high-level summary of how frequently each justice-oriented goal appears across the dataset, offering a complementary quantitative view of the prevalence and distribution of reparative responses. In our analysis of 1,060 AI incidents, 54% involved actions related to Acknowledgment and 47% involved Attribution. By contrast, only 25% of cases included any form of Remedy, and 20% showed evidence of Reform. Figure 1 presents the distribution of specific actions across these goals. The most frequent action was communication from perpetrators, appearing in 541 cases (51%). In comparison, more impactful forms of follow-through were rare: only 7% of cases involved legal or policy reform, 6% mentioned compensation, and just 4% resulted in the discontinuation of a product or feature.
@@
Attribution Investigation of Harm Regulators, Media,
AI Corporations
Internal corporate audits; Government audits; Investigative journalism
Initiation of Legal Actions Affected Users, Advocacy Groups, Regulators Pre-action letter; Initiation of lawsuits
Remedy Repercussions Regulators,
AI Corporations
Fines; Criminal charges; Dismissals
Compensation AI Corporations,
End-Users
Refunds; Settlements
Removal of Content
or Product Recall
AI Corporations Partial dataset removal; Product safety recall
Reform Discontinuation AI Corporations Permanent product or service shutdown
AI Design Changes AI Corporations Algorithm redesign; Adjustments to product safety protocols
Law or Policy Change Regulators,
AI Corporations
Public legislation updates; Internal policy updates
Table 1: Taxonomy of Reparative Actions. This table presents our taxonomy of reparative actions, organized by four overarching goals: Acknowledgment, Attribution, Remedy, and Reform. For each goal, we list corresponding action types, the major stakeholders involved, and common forms these actions take.